OpenStudy (anonymous):

1. Rewrite with only sin x and cos x. sin 3x - cos x

OpenStudy (anonymous):

sin3x-cosx =sin(2x+x)-cosx =sin2x*cosx+sinx*cos2x-cosx =2sinx*cos^2(x)+sinx*(cos^2(x)-sin^2(x))-cosx =3sinx*cos^2(x)-sin^3(x)-cosx

OpenStudy (anonymous):

Using the following formulas sin(a+b)=sina*sinb+cosa*cosb sin2x=2sinx*cosx cos2x=cos^2(x)-sin^2(x)